OpenStudy (anonymous):

the mean of 100 numerical observations is 47.83 what is the value of the sum of all 100 numbers

OpenStudy (john_es):

The mean is defined by, \[\overline{x}=\frac{x_1+x_2+x_3+x_3+\ldots+x_N}{N}=\frac{S}{N}\]They ask you for the number S, provided N=100, and the mean is 47.83.

OpenStudy (anonymous):

thanks